OpenStudy (anonymous):

A landscaper is designing a flower garden in the shape of a trapezoid. She wants the length of the shorter base to be 3 yards greater than the height, and the length of the longer base to be 5 yards greater than the height. For what height will the garden have an area of 360 square yards? Round to the nearest tenth of a yard.

OpenStudy (anonymous):

if you assume that the height is x; then the short base is x+3;; and the long base is x+5;

OpenStudy (anonymous):



OpenStudy (anonymous):

the area of trapezoid is ( short base + long base) *hieght /2 = (x+3 + x+5)*x/2 = 360;

OpenStudy (anonymous):

\[x^2 + 4x -360 ;\]

OpenStudy (anonymous):

to find root of the equation you have to use \[x1 = (-b +\sqrt{b^2 -4ac})/2 \] \[x2 = (-b -\sqrt{b^2 -4ac})/2 \]

OpenStudy (anonymous):

b= 4 and c= -360 and a= 1; after substitue in the formula you will get x1 = 17.075 after rounding you will get 17.1- this the answer x2 = -21.075 ( but the height is nutral number !!!!) this is not a solution the answer is 17.1